OpenStudy (anonymous):

PLEASE HELP! IM STUCK! Simplify log6 4x + log6 5x + log6 2x

OpenStudy (anonymous):

\[\log _{6}4x+\log _{6}5x + \log _{6}2x\]\[\log_{6}(4x \times5x \times2x)\]\[\log _{6}(120x ^{3})\]

OpenStudy (anonymous):

I dont understand! :(

OpenStudy (anonymous):

log6(4x times 5x times 2s)

OpenStudy (anonymous):

A property of log includes \[\log _{a}B+\log _{a}C = \log _{a} (B C)\]

OpenStudy (anonymous):

er, 2x. typo.